# Cryptographic Censorship

### 05 Feb 2024 v1

@@ 4.2 Accommodating gravitational pseudorandomness
We will now show that singularities with $a<d-1$ are compatible with gravitational pseudorandomness. Consider an asymptotically AdS spacetime with a timelike naked singularity $\Gamma=\Gamma_{n(\mathscr{I})}$. We will work with large (but possibly finite) $N$ and demand only that effective field theory as defined by a bulk UV cutoff be valid in the asymptotic region. At finite $N$ we will consider the failure of EFT due to large UV corrections as the diagnostic of a singularity. [^38] Here we are taking EFT to be resulting solely from a UV cutoff rather than including the more complete version of an EFT proposed by [31], which also includes a complexity restriction.
@@
Figure 13: From the low-energy perspective, there is an unknown, UV-sensitive boundary condition assigned at the singularity; for example, modes that fall into the singularity are turned into a set of outgoing modes with random phases $\theta_{ij}$.
From the low-energy perspective, there is a boundary condition assigned at the singularity whose precise nature is unknown. [^39] Since quantum gravity is expected to lack global symmetries [68, 69, 70], any approximate symmetries emergent in the semiclassical EFT should be badly violated by this boundary condition. Now, consider a set of ingoing wavepackets |$\ket{i}_{\rm in}$ that eventually fall into the singularity. We might expect that these modes scatter off the singularity in an apparently random manner. One toy model of such dynamics turns ingoing modes into outgoing modes of seemingly random phases $\theta_{ij}$, schematically,
|$\ket{i}_{\rm in}\overset{\delta t}{\rightarrow}\frac{1}{\sqrt{\mathcal{N}}}\sum_{j}e^{i\theta_{ij}}\ket{j}_{\rm out},$ (4.4)
where $\mathcal{N}$ is a normalization constant. The phases $\theta_{ij}$ could in principle correspond to pseudorandom phases, which would be a consequence of pseudorandom dynamics in the boundary description [^40] There are constraints from overall conservation of ADM energy and gauge charge. But there will be many collections of wavepackets that have the same energy, so that (4.4) would be consistent with charge conservation. (though of course determining what $\theta_{ij}$ actually is requires access to UV physics). See, e.g., [33] for explicit constructions of pseudorandom states using pseudorandom phases. Our approach instead will be to show that once the singularity is of classical or semi-Planckian size, then there are enough EFT modes in a code subspace in which all of the states have the same leading order geometry to accommodate gravitational pseudorandomness. On the other hand, if the singularity is Planckian, we show that we do not expect this. As our analysis is asymptotic, we do not require a sharply defined geometry near the singularity.
@@ 5 Discussion
Since information escapes from the black hole interior under inclusion of quantum corrections, there is no obvious definition of a quantum event horizon (since all information eventually reaches $\mathscr{I}$ and causal structures are only emergent in the classical limit). It is tempting to use our results to define quantum event horizons — and by extension quantum black holes (i.e., at finite $N$): states undergoing gravitationally pseudorandom time evolution with a semiclassical bulk dual in the large-$N$ limit. A definition, however, would require a proof of a converse to Cryptographic Censorship: a statement that CFT states with a semiclassical emergent bulk featuring a horizon typically exhibit gravitational pseudorandomness. A proof of such a converse would additionally facilitate a complete resolution of the teleology problem in AdS. The tentative link between chaos, thermalization, and pseudolandomness [76, 77, 34, 65, 78, 79] could then potentially be used to understand the horizon thermodynamics [^48] As opposed to the thermodynamics of the apparent horizon, which are by now well-understood [54, 28, 80]. of highly dynamical black holes and the statistical interpretation of the Hawking area theorem, which remains elusive [27, 81].
Implications for the Bulk
There is an immediate implication of Cryptographic Censorship for the relationship between complexity and black hole formation: bulk processes that source a large amount of pseudorandomness (e.g., a bulk quantum computer with the power to do gravitationally pseudorandom computations) must typically collapse themselves into black holes. Prior descriptions of black hole formation typically involve energetics; to our knowledge, this is the first proof establishing a causal relation between bulk quantum computations and black hole formation. Interestingly, this appears to bear some relation to the proposal of [31] that effective field theory requires both an energy and a complexity cutoff: a bulk process sourcing large amounts of pseudorandomness must result in high reconstruction complexity. While pseudorandomness is itself not particularly complex to implement, reconstruction of pseudorandom quantities is by definition of high complexity class. In this case, Cryptographic Censorship may act to protect a computationally bounded asymptotic observer from detecting large violations of effective field theory due to computational complexity by forcing formation of an event horizon, depending on where the event horizon forms.
@@
    Our definition of the size of a singularity admits a middle class of singularities that disappear in the strict larle-$N$ limit but which could accommodate gravitational pseudorandomness in code subspaces constructed from $O(1)$ time bands. It would be interesting to understand the significance, if any, and frequency of semi-Planckian naked singularity portions.
@@ Acknowledgements
It is a pleasure to thank Jordan Cotler, Jordan Docter, Roberto Emparan, Daniel Harlow, Gary Horowitz, Hsin-Yuan (Robert) Huang, Andrea Puhm, Arvin Shahbazi-Moghaddam, Natalie Paquette, Jon Sorce, Edward Witten, and Michelle Xu for valuable discussions. This work is supported in part by the Department of Energy under Early Career Award DE-SC0021886 (NE, ÅF, and LY) and QuantISED DE-SC0020360 contract no. 578218 (NE), by the Heising-Simons Foundation under grant no. 2023-4430 (NE and AL), the Sloan Foundation (NE), the Templeton Foundation via the Black Hole Initiative (NE and EV), the Gordon and Betty Moore Foundation (EV), the Packard Foundation (AL), and the MIT department of physics (NE). The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton or Moore Foundations.
@@ Appendix A Proof of Theorem 3
Note that by the assumptions on the sequences $\{U^{(N)}\ket{\psi^{(N)}}\}_{N}$ and $\{\widehat{U}^{(N)}\ket{\psi^{(N)}}\}_{N}$, for every sufficiently large $N$, $U^{(N)}\ket{\psi^{(N)}}$ and $\widehat{U}^{(N)}\ket{\psi^{(N)}}$ satisfy 1. and 2. in Theorem 3.

### 19 Mar 2024 v2

@@ 4.2 Accommodating gravitational pseudorandomness
We will now show that singularities with $a<d-1$ are compatible with gravitational pseudorandomness. Consider an asymptotically AdS spacetime with a timelike naked singularity $\Gamma=\Gamma_{n(\mathscr{I})}$. We will work with large (but possibly finite) $N$ and demand only that effective field theory as defined by a bulk UV cutoff be valid in the asymptotic region. At finite $N$ we will consider the failure of EFT due to large UV corrections as the diagnostic of a singularity.
@@
Figure 14: From the low-energy perspective, there is an unknown, UV-sensitive boundary condition assigned at the singularity.
From the low-energy perspective, there is a boundary condition assigned at the singularity whose precise nature is unknown. [^38] Since quantum gravity is expected to lack global symmetries [69, 70, 71], any approximate symmetries emergent in the semiclassical EFT should be badly violated by this boundary condition.
In what follows, we will be agnostic about the UV-sensitive boundary condition that implements pseudorandomness in the fundamental theory. Our approach instead will be to show that once the singularity is of classical or semi-Planckian size, then there are enough EFT modes in a code subspace in which all of the states have the same leading order geometry to accommodate gravitational pseudorandomness. On the other hand, if the singularity is Planckian, we show that we do not expect this. As our analysis is asymptotic, we do not require a sharply defined geometry near the singularity.
@@ 5 Discussion
Since information escapes from the black hole interior under inclusion of quantum corrections, there is no obvious definition of a quantum event horizon (since all information eventually reaches $\mathscr{I}$ and causal structures are only emergent in the classical limit). It is tempting to use our results to define quantum event horizons — and by extension quantum black holes (i.e., at finite $N$): states undergoing gravitationally pseudorandom time evolution with a semiclassical bulk dual in the large-$N$ limit. A definition, however, would require a proof of a converse to Cryptographic Censorship: a statement that CFT states with a semiclassical emergent bulk featuring a horizon typically exhibit gravitational pseudorandomness. A proof of such a converse would additionally facilitate a complete resolution of the teleology problem in AdS. The tentative link between chaos, thermalization, and pseudorandomness [77, 78, 34, 66, 79, 80] could then potentially be used to understand the horizon thermodynamics [^46] As opposed to the thermodynamics of the apparent horizon, which are by now well-understood [55, 28, 81]. of highly dynamical black holes and the statistical interpretation of the Hawking area theorem, which remains elusive [27, 82].
Gravitational Pseudorandomness from a Bulk Quantum Computer?
The reader might wonder: what if we put a quantum computer in the bulk, generating a large amount of pseudorandomness directly from dynamics of the bulk fields? Must this incur a horizon? If the quantum computer does not incur violations of EFT in the causal wedge, then our theorem applies and indeed, a gravitationally pseudorandom computer must incur a horizon. If the quantum computer generates gravitational pseudorandomness in such a way that the causal wedge is no longer described by EFT — e.g., if geometric invariants such as the volume scale inversely with $G_{N}$ as in the “quantum volatile” spacetimes of [59] — then the assumptions of our theorem are violated, namely that the $N\rightarrow\infty$ limit is described by QFT in curved space, so a horizon is not required to form.
It is interesting that bulk processes that source a large amount of pseudorandomness (e.g., a bulk quantum computer with the power to do gravitationally pseudorandom computations) and do not violate EFT in the causal wedge must typically collapse into black holes. Prior descriptions of black hole formation typically involve energetics; to our knowledge, this is the first proof establishing a causal relation between bulk quantum computations and black hole formation (though see, e.g., [83] for a discussion of constraints on bulk quantum computations behind horizons). Interestingly, this appears to bear some relation to the proposal of [31] that effective field theory requires both an energy and a complexity cutoff: a bulk process sourcing large amounts of pseudorandomness must result in high reconstruction complexity. While pseudorandomness is itself not particularly complex to implement, reconstruction of pseudorandom quantities is by definition of high complexity class. In this case, Cryptographic Censorship may act to protect a computationally bounded asymptotic observer from detecting large violations of effective field theory due to computational complexity by forcing formation of an event horizon, depending on where the event horizon forms.
@@
    Our definition of the size of a singularity admits a middle class of singularities that disappear in the strict large-$N$ limit but which could accommodate gravitational pseudorandomness in code subspaces constructed from $O(1)$ time bands. It would be interesting to understand the significance, if any, and frequency of semi-Planckian naked singularity portions.
@@ Acknowledgements
It is a pleasure to thank Chris Akers, Jordan Cotler, Jordan Docter, Roberto Emparan, Daniel Harlow, Gary Horowitz, Hsin-Yuan (Robert) Huang, Andrea Puhm, Juan Maldacena, Arvin Shahbazi-Moghaddam, Natalie Paquette, Jon Sorce, Edward Witten, and Michelle Xu for valuable discussions. This work is supported in part by the Department of Energy under Early Career Award DE-SC0021886 (NE, ÅF, and LY) and QuantISED DE-SC0020360 contract no. 578218 (NE), by the Heising-Simons Foundation under grant no. 2023-4430 (NE and AL), the Sloan Foundation (NE), the Templeton Foundation via the Black Hole Initiative (NE and EV), the Gordon and Betty Moore Foundation (EV), the Packard Foundation (AL), and the MIT department of physics (NE). The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton or Moore Foundations.
@@ Appendix A Proof of Theorem 3
By the assumptions on the sequences $\{U^{(N)}\ket{\psi^{(N)}}\}_{N}$ and $\{\widehat{U}^{(N)}\ket{\psi^{(N)}}\}_{N}$, for every sufficiently large $N$, $U^{(N)}\ket{\psi^{(N)}}$ and $\widehat{U}^{(N)}\ket{\psi^{(N)}}$ satisfy 1. and 2. in Theorem 3.